OpenStudy (anonymous):

Find the derivative of f(x) = 8x + 4 at x = 9.

OpenStudy (anonymous):

what is the derivative of 8x+4?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@dan815 @ganeshie8 @Miss.SweetiePie

OpenStudy (welshfella):

if y = ax^n derivative is anx^(n-1)

OpenStudy (welshfella):

now plug in a = 8 and n = 1 into ax^n and see what you get

OpenStudy (anonymous):

8x^1

OpenStudy (welshfella):

ok now plug those values of a and n into anx^(n-1) to get the derivative

OpenStudy (anonymous):

these r the options 0 4 8 9

OpenStudy (welshfella):

derivative = 8*1x^(1-1) = ?

OpenStudy (anonymous):

thanks @stratemeyerjr

OpenStudy (welshfella):

= 8 x^0 what is x^0 equal to ?

OpenStudy (anonymous):

thank @welshfella :) he put in the effort